If 2 cards are selected from a standard deck of 52 cards without replacement, find these probabilities.
a. Both are spades.
b. Both are the same suit.

c. Both are kings.

Answer by jim_thompson5910(35256)

I'll do the first two to get you started.

a)

There are 13 spades out of 52 cards total. So P(drawing spade on first draw) = 13/52

After the first draw, there are 12 spades out of 51 total. So P(drawing spade on second draw) = 12/51. Remember, we're NOT replacing the cards.


So P(Both Spades) = P(drawing spade on first draw)*P(drawing spade on second draw)=(13/52)(12/51) = 156/2652 = 1/17


So the probability is 1/17
--------------------------------------------------------------------
b)

Regardless of what suit is chosen on the first draw, we'll have 12 cards left over in that selected suit (since all 4 suits have 13 cards) out of 51 total


So P(Both same suit) = 12/51


Note: the first card does not matter because the second card determines the entire probability. If you're still unsure, do some simulations (ie actually pull out a deck of cards and start drawing). Say on the first draw you pick a heart (actual face value doesn't matter). You'll then have 12 heart cards left out of 51 cards total.


--------------------------------------------------------------------
c)

I leave this one for you to try. Let me know if you still need help.